OpenStudy (anonymous):

Which is the correct algebraic equation for the sentence? Fifty-two equals two more than a number multiplied by 5.

OpenStudy (anonymous):

52 = 2 + 5x

OpenStudy (agreene):

\[52=2+5\phi \]

OpenStudy (anonymous):

options A 52 = 2 + 5 + s B 52 = 2 - 5s C 52 = 2 + 5s D 52 = 5s - 2

OpenStudy (anonymous):

c is the answer right??????

OpenStudy (agreene):

C. is the same was what me and @sethw wrote.